OpenStudy (anonymous):

1/2-2/6 and i need shown work

OpenStudy (anonymous):

get a common denominator. the first part is to look at 2 and 6. since you can get 6 by multiplying 2 by 3 you can work it so take 1/2 and multiply by 3/3 which is 1 which doesn't change the fraction at all. 1/2*3/3 = 3/6 then 3/6 - 2/6 = 1/6

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

if you dont understand any part i can re-explain. see im not all douche.

OpenStudy (anonymous):

3/4+2/3

OpenStudy (anonymous):

and u can be

OpenStudy (anonymous):

this one requires a little more work. you have to figure the common denominator which will be the least common multiple between 3 and 4. if you notice that 3*4 = 12 and then see that no other number less than 12 is divisible by BOTH 3 and 4, you have found it. (if you use a bigger number than 12 divisible by both 3 and 4, it is ok since you can factor out at the end!) so then 3/4 * 3/3 = 9/12 and 2/3 *4/4 = 8/12 so 8/12 + 9/12 = 17/12 you cannot factor anymore so this is the answer. unless the problem says you cant have improper fractions then this is ok

OpenStudy (anonymous):

how old r u

OpenStudy (anonymous):

20

OpenStudy (anonymous):

wow and u used to argue with me this may be the beginning of a betiful friendship

OpenStudy (anonymous):

11/12-1/6

OpenStudy (anonymous):

dude i'll help anyone with math

OpenStudy (anonymous):

what did i say about calling me dude u short stump of a person

OpenStudy (anonymous):

lmao dude im not short at all

OpenStudy (anonymous):

just cause im 11 doesnt mean u dont gotta show me respect ill bring u down to your kneews wen i see u if i do

OpenStudy (anonymous):

ok so its the same thing a commone denominator between 12 and 6. well you know that 6*2 = 12 so this works we have 1/6 *2/2 = 2/12 so 11/12-2/12 = 9/12 now notice you can factor the numerator and denominator to get 3*3 / 3*4 and cancel a set of 3's and we are left with 3/4

OpenStudy (anonymous):

haha

OpenStudy (anonymous):

IM NOT JOKING U NEED TO TALK TO ME LIKE YOUVE GOT SENSE /5/6+3/4

OpenStudy (anonymous):

haha why dont you try this one out to see if you can do it? go through it exactly the way i did

OpenStudy (anonymous):

if i could do it i wouldnt be askin for help now would i

OpenStudy (anonymous):

im done i did it

OpenStudy (anonymous):

haha good job buddy what'd you get for the answer?

OpenStudy (anonymous):

the answer is 18 5/6

OpenStudy (anonymous):

mother flutterer dont call me buddy

OpenStudy (anonymous):

you are wrong dude

OpenStudy (anonymous):

5/6 + 3/4

OpenStudy (anonymous):

common denominator is 12. 5/6*2/2 = 10/12 3/4*3/3 = 9/12 10/12 + 9/12 = 19/12

OpenStudy (anonymous):

no improper fraction its actually 1 7/12

OpenStudy (anonymous):

yeah if its mixed fractions

OpenStudy (anonymous):

r u going from me to other people/1/2+7/8

OpenStudy (anonymous):

haha maybe ok so common denominator between 2 and 8 is 8 so 2*4 = 8 so 1/2 * 4/4 = 4/8 then 4/8 + 7/8 = 11/8 or 1 3/8